evaluate ∫ cos 2x + 2 sin^2 x / sin^2 x dx

evaluate  ∫ cos 2x + 2 sin^2 x / sin^2 x  dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{cos2x + 2sin^{2}x}{sin^{2}x}dx
cos2x = 1 - 2sin^{2}x
cos2x + 2sin^{2}x = 1
I = \int \frac{1}{sin^{2}x}dx
I = \int cosec^{2}xdx
I = -cotx + constant